OpenStudy (anonymous):

Given the equation 5x+10= -30, solve for x and choose the correct solution. k minus seventeen equals negative fourteen Three k minus two and one third equals nine Six k minus two equals sixteen Seven plus two k equals twenty

OpenStudy (anonymous):

@agb

OpenStudy (anonymous):

http://openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

5x+10=-30 5x=-30-10 5x=-40 x=-40/5 x=-8

OpenStudy (anonymous):

thanks @n1742

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

k-17=-14 k=-14+17 k=3

OpenStudy (anonymous):

6k-2=16 6k=16+2 6k=18 k=18/6 k=3

OpenStudy (anonymous):

7+2k=20 2k=20-7 2k=13 k=13/2

OpenStudy (anonymous):

@huntin14